# First-Order Vortex Lattice Melting in Bilayer Ice: A Monte Carlo Method Study

### v1

@@ II.2.2 The measured quantities
and the four phases, along with their boundaries expressed in terms of the parameters $a$, $b$ , $c$ and $\Delta$, are listed in the following table:
@@
One may wonder why the two-layer 6V model is the same as the single-layer square lattice 6V model. Let us now briefly analyze a few locations of the phase transition boundary. The first point is $(a/c,b/c)=(1,0)$ which is the point of phase transition between the AFM and FM phases. As illustrated, the FM-phase is full of vertices with weight $a$ and the AF-phase is full of vertices with weight $c$. At the point of phase transition of the two phases, i.e., phase AFM and phase FM, free energies are equal defined as following,
@@
$S_{\text{FM}}$ $=k_{B}\text{ln}2,$ (18)
$S_{\text{AFM}}$ $=k_{B}\text{ln}2.$
Along the $b=0$ axis, the system in the FM phase only has either all a1-type vertices or all a2-type vertices. In other words, there are two microscopic states. Similarly, in the AFM phase, there exist configurations with only all c1 or all c2 type vertices. Therefore $a/c=1$ is the critical point satisfying Eq. 17 at $b=0$.
Figure 5: Global phase diagram in plane ($a/c-b/c$) with $V=1$ and snapshots. (a) phase diagram containing AFM, FM and disorder phases (b) the typical snapshot in the FM phase. The AFM-D transition is of Ising type. (c)Two configurations of the AFM phase for the Z2 symmetry breaking.
III.2 Ising type not BKT type of transiton along $a=b$
The phase transitions of the 6V model is identical on both the square lattice and double-layer honeycomb lattice. For example, along $a=b$, Ref. [25] confirms the critical point is at $\beta_{c}=\text{ln}(2)$. The critical point locates at
@@ III.2 Ising type not BKT type of transiton along $a=b$
In our MC simulation, as shown in Fig. 6 (a)-(b), he intersection of $m_{x}^{-}$ and its Binder ratio $Q$ locates at $a_{c}=b_{c}=1/2$.
However, the type of the phase transition changes. In the $F$ model, the phase transition belongs bo be the BKT type, as evidenced by the non-divergent behavior of specific heats in Fig. 4 in Ref. [25]. In our bilayer 6V model, the type should be Ising type.
The first signature is the value the Ising type critical exponent $\nu=1/y_{t}=1$ satisfys finite size scaling equation defined as follows
@@
Using the data clapse method, $Q$ is plotted as function of $(a-a_{c})L^{y_{t}}$, data of different sizes overlap as shown in Fig. 6 (c). This phenomenon supports the conclusion that the observed phase transition is indeed the Ising type.
This type of Ising phase transition can also be analyzed from the perspective of symmetry breaking. In the AFM phase, as depicted in Fig. 5(c), there is a twofold degeneracy in configuration. In the horizontal direction, one configuration is characterized by a c1-c2-c1-c2 vertex arrangement, while the other features a c2-c1-c2-c1 arrangement. Moreover, the relationship between these two configurations is achieved by flipping the states of all legs. This means that from disorder to AFM, there is a Z2 symmetry breaking leading to the Ising transition.
To further confirm our code, we also simulate the F model on the one-layer square lattice, the $C_{V}$ indeed does not diverge at $a_{c}$ as shown in Fig. 6 (d), consistent with the result in Ref. [25].
Figure 7: Global phase diagram with $V\neq 1$ and snapshots. (a) A schematic phase diagram (b) the AF phase (c) 2/3 VL phase (d) 1/2 VL phase. In fact, 1/2 VL is accompanied by $x$-direction ferromagnetic order and 1/3 VL is accompanied by AFM order, see appendix A.
IV phase diagram with $V\neq 1$
In this section, we introduce a non-zero value for $v$ in Eq. 1 to discuss the effect the vortex excitation.
As in Ref. [29], it is experimentally possible to manually insert of delete vortices [46, 47], despite the fact that the factors are added manually here.
This phase diagram is shown in the $a-V$ plane, where $V=e^{-v}$, while simultaneously maintaining a fixed cut along the cut $b=a$.
@@
Figure 7 (a) shows the global pahse diagram, which contains the AFM, disorder, 1/2 and 2/3 VL phases. When the vortex excitation is considered, the phase diagram becomes rich. In Fig. 7 (c), the configuration of the 2/3-VL phase is shown. The dual lattice of the honeycomb lattice is the triangular lattice. In this phase, two of the three sets of sublattices within the triangular lattice are occupied by positive and negative vortices, respectively, while the remaining set of sublattices contains no vortices. In Fig. 7 (d), the configuration of the 1/2-VL phase is shown. In this phase, positive and negative vortices alternate sequentially along the $y$-direction, forming connected stripes.
Here, we analyze the distribution of different phases within the phase diagram. When $a$ and $b$ are very small, $a=b\ll c$ and $V\ll c$, the system resides in the AFM phase, represented by the dark red region. The upper right region, $a=b\gg c$ and $V\gg c$, the system corresponds to the 1/2 VL phase. The system has almost no type $c$ vertices. At the same time $V$ drives the system to form vortices consisting of vertices of type $a$, $b$ only, as shown in Fig. A2 (a).
@@
[…]
@@
In fact, 1/2 VL is accompanied by $x$-direction ferromagnetic order and 1/3 VL is accompanied by AFM order, see appendix A.
@@ IV.2 Vortex structure factors
To further understand the spin vortex lattice phase, the the structure factor in $\bm{q}$-space is introduced as
@@
where $n_{i}$, $n_{j}$ represents 1, 0, in the face of the honeycomb lattices, i.e., triangular lattice. The symbols $\bm{r_{i}}$ and $\bm{r_{j}}$ are the center coordinates of the vortex. In real space, if the density obeys configurations of the form ($101010\cdots$) or ($010101\cdots$), the wave vector corresponding to the maximum value of $S(\bm{q})$ should be located at $\mathbf{q}=(\pi,0)$ [48].
@@
In Fig. 8 (a), for the 2/3 VL phase, $S(q_{x},q_{y})$ is obtained by using Eq. 21 with a lattice size $L=12$. One of the brightest point is located at
@@
pointed by the red arrow. The position of the peaks reflects the translational symmetry of the vortex lattice. Assuming that the side length of the honeycomb lattice is 1, the spacing between the two pink vortices $\bm{r_{i}}$ and $\bm{r_{j}}$ is
@@
The positon in Eq. 22 can be obtained by using
@@
For the 1/2 VL lattice, $S(q_{x},q_{y})$ is shown in Fig. 8 (b). The brightest point (by the blue arrow) is located at
@@
and the second brightest point (by the red arrow) is located at
@@
This is because in real space the translation vectors are respectively $\Delta x=2\sqrt{3}$ and $\Delta y=3$.
In total, we have identified two VL phases in the bilayer 6V model with vortex weight. These ordered states can be distinguished by examining either their respective structure factors or the relevant order parameters.
@@ IV.3 Size effects from periodic boundaries
Figure 9 shows the size effects of the vortex density in a 2/3 VL with $L=4,6,8,10$, respectively. The honeycomb lattice is a complex lattice with two sites in its smallest cell, hence having an even number of lattice sites in the horizontal direction, i.e., $L\%2==0$.
@@
Table 3: The table shows the maximum theoretical value of the vortex density $\rho_{v}$ as a function of system size $L$
@@
where the symbol $\lfloor$.$\rfloor$ . ⌋ denotes rounding down. We first explain Eq. 28 using $L=6$, two sites labeled as 1, 2 in horizontal direction, totally $L/2$ bricks in each row as shown in denominator.
@@
Through the analysis above and the densities presented in Table 3, in the 2/3 VL phase, the density remains constant for sizes where $L\%6==0$. Therefore, when observing physical quantities later, we only simulate systems whose size is a multiple of 6.
@@ IV.4 Detailed transitions between the phases
IV.4.1 The Ising transition between the AFM and disorder phases at $V<1$
@@
In Figs. 10 (a) and (b), the specific heat $C_{V}$ and $C_{vx}$ exhibit divergence with respect to sizes during the phase transition between the AFM and disorder phases. This behavior stands in contrast to the convergence of $C_{V}$ in the single-layer 6V $F$ model [25], as shown in Fig. 6 (d).
@@
The possible reasons are as follows: $V<1$ implies the absence of the disordered phase as depicted by the absence of vortices in Fig. A2 (c). Additionally, the AFM phase has no vertex, as illustrated in Fig. 7 (d). Consequently, the phase transition between these two phases does not involve vortices. Moreover, the transition from simple AFM ordering to disorder involves $Z2$ symmetry breaking.
Figure 10: The Ising transition between the AFM and Disorder phases at $V<1$, $a=b$ and $V=0.5$. (a) $C_{V}$ (b) $C_{vx}$, (c) $Q$ (d) data collapse of $Q$, using $y_{t}=1$. The system size ranges from 8 to 24.
IV.4.2 The transitions between the D and 1/2 VL phases at $V>1$
The phase transition between the disorder and 1/2 VL phase is discussed. By fixing the parameter $V=4$, and scanning $a$ and keeping $a=b$, the vortex density $\rho_{vx}$, and $C_{V}$ are measured for different sizes $L=16-48$.
As shown in Fig 11 (a) and (b), the jumps of $\rho_{vx}$ and peaks of $C_{V}$ show the signatures of phase transitions. The divergence of $C_{V}$ confirms the phase transition is not the BKT type [29]. In Figs. 11 (c)-(d), the histogram of $E/N$ and $m_{s}$ are obvious double-peaked at the parameters $a=b=2$. This indicates the phase transition is first-order type.
The signature of the first-order transition can also obtained by fitting the critical exponent. Using the package for finite size scaling [49], $\nu$ is obtained as $0.48(8)$ using Eq. 20.
The data for sizes $L=32,36,40$ overlaps very well. The scaling dimension $y_{t}=1/0.48=2.08\approx 2$ equals to the system dimension $d=2$ when the first-order transition occurs [50].
Figure 11: The signatures of the first-order phase transition at $V=4$. The quantities (a) $\rho_{vx}$ (b) $C_{V}$ with system sizes $L=16-48$. The double-peaked distributions (c) $P(E/N)$ and (d) $P(m_{s})$ at $a=2.004(1)$. (e) $Q_{s}$ and (f) data collapse of $Q$ and obtaining $\nu=0.48(8)$ [49].
IV.4.3 The transition between the D and 2/3 VL phases
Figure 12: The signatures of the first-order phase transition. The double-peaked distribution of (a) $P(E/N)$ and (b) $P(\rho_{vx})$ at $V=7$ and $a=0.35$. (c) $C_{V}$ versus $a$, (d) $a_{c}$ versus $1/L$. (e) $E/N$ versus MC steps and the convergent time is about $10^{6}$.
The transition from the 2/3 VL phase to the disorder phase is also first order. In Figs. 12(a) and (b), the double peaks in the distribution of $P(E)$ and $P(\rho_{vx})$ indicating a first-order phase transition.
Different from 1/2 VL phase as shown in Fig. 11, this phase transition has an obvious size effect. Figure 12(c) illustrates that the specific heat peak shifts to the right as the size increases. Fortunately, through finite-size scaling defined as follows
@@ IV.4.3 The transition between the D and 2/3 VL phases
The error bar (2) is calculated by the following equations. First, if one fits $y=kx+b$, then the standard deviation of the intercept b is
@@
where $n$ is the number of points involved, $s_{y}$ is the standard deviation of the observation $y_{i}$ and can be expressed as
@@
where $n-2$ is the degree of freedom. If one fits the data using software “gnuplot”, the result is $a_{c}\left(\infty\right)=0.476(5)$ consistent with the above result within the errorbar.
To show that the results are convergent, Fig. 12(e) displays $E/N$ versus MC steps. The converging time is about $10^{6}$ steps. The converged energy is approximately 0.12, denoted as $e_{c}$. $e_{c}$ is also marked by a dashed line in Fig. 12(a), sitting between the two peaks.

### v2

@@ III.1 Global phase diagram with $\bf V=1$
and the four phases, along with their boundaries expressed in terms of the parameters $a$, $b$ , $c$ and $\Delta$, are listed in Table  2.
@@
One may wonder why the two-layer 6V model is the same as the single-layer square lattice 6V model. Let us now briefly analyze a few locations of the phase transition boundary. The first point is $(a/c,b/c)=(1,0)$ which is the point of phase transition between the AFM and FM phases. As illustrated in Fig. 5 (a), the FM-phase is full of vertices with weight $a$ and the AF-phase is full of vertices with weight $c$. At the point of phase transition of the two phases, i.e., phase AFM and phase FM, free energies are equal and defined by the following:
@@
$S_{\text{FM}}$ $=S_{\text{AFM}}=k_{B}\text{ln}2.$ (18)
Along the $b=0$ axis, the system in the FM phase only has either all a1 type vertices or all a2 type vertices. In other words, there are two microscopic states. Similarly, in the AFM phase, there exist configurations with only all c1 or all c2 type vertices. Therefore $a/c=1$ is the critical point satisfying Eq. 17 at $b=0$.
Figure 5: Global phase diagram in the plane ($a/c-b/c$) with $V=1$ and snapshots. (a) phase diagram containing AFM, FM and disordered phases (b) the typical snapshot in the FM phase. The AFM-D transition is of Ising type. (c) Two configurations of the AFM phase for the Z2 symmetry breaking.
III.2 Ising type not BKT type in transition along $a=b$
The phase transitions of the 6V model are identical on both the square lattice and the double-layer honeycomb lattice. For example, along $a=b$, Ref. [26] confirms the critical point is at $\beta_{c}=\text{ln}(2)$. The critical point locates at:
@@ III.2 Ising type not BKT type in transition along $a=b$
In our MC simulation, as shown in Fig. 6 (a)-(b), the intersection of $m_{x}^{-}$ and its Binder ratio $Q$ locates at $a_{c}=b_{c}=1/2$.
However, the type of the phase transition changes. In the $F$ model, the phase transition belongs to be the BKT type, as evidenced by the non-divergent behavior of specific heats in Fig. 4 in Ref. [26]. In our bilayer 6V model, the type should be of Ising type.
The first signal is that, for the Ising phase transition, the critical exponent $y_{t}=1$. Using the finite size scaling equation defined as follows:
@@
and the data [ collapse] method, $Q$ is plotted as function of $(a-a_{c})L^{y_{t}}$, and data of different sizes overlap as shown in Fig. 6 (c). This phenomenon supports the conclusion that the observed phase transition is indeed the Ising type.
This type of Ising phase transition can also be analyzed from the perspective of symmetry breaking. In the AFM phase, as depicted in Fig. 5 (c), there is a twofold degeneracy in the configuration. In the horizontal direction, one configuration is characterized by a c1-c2-c1-c2 vertex arrangement, while the other features a c2-c1-c2-c1 arrangement. Moreover, the relationship between these two configurations is achieved by flipping the states of all legs. This means that from disorder to AFM, there is a Z2 symmetry breaking leading to the Ising transition.
To further vindicate our code, we also simulate the F model on the one-layer square lattice; the $C_{V}$ indeed does not diverge at $a_{c}$ as shown in Fig. 6 (d), consistent with the result in Ref. [26].
Figure 7: Global phase diagram with $V\neq 1$ and snapshots. (a) A schematic phase diagram (b) the AF phase (c) 2/3 VL phase (d) 1/2 VL phase. In fact, 1/2 VL is accompanied by a $x$-direction ferromagnetic order and 1/3 VL is accompanied by an AFM order (see appendix A).
IV phase diagram with $\bf V\neq 1$
In this section, we introduce a non-zero value for $v$ in Eq. 1 to discuss the effects of vortex excitation.
As in Ref. [30], it is experimentally possible to manually insert or delete vortices [47, 48], despite the fact that the factors are added manually here.
This phase diagram is shown in the $a-V$ plane, where $V=e^{-v}$, while simultaneously maintaining a fixed cut along $b=a$.
@@ IV.1 Global phase diagram
Figure 7 (a) shows the global phase diagram, which contains the AFM, disordered, 1/2 and 2/3 VL phases. When the vortex excitation is considered, the phase diagram becomes rich in features. In Fig. 7 (c), the configuration of the 2/3-VL phase is shown. The dual lattice of the honeycomb lattice is the triangular lattice. In this phase, two of the three sets of sublattices within the triangular lattice are occupied by positive and negative vortices, respectively, while the remaining set of sublattices contains no vortices. In Fig. 7 (d), the configuration of the 1/2-VL phase is shown. In this phase, positive and negative vortices alternate sequentially along the $y$-direction, forming connected stripes.
Here, we analyze the distribution of different phases within the phase diagram. When $a$ and $b$ are very small, $a=b\ll c$ and $V\ll c$, the system resides in the AFM phase, represented by the dark red region. In the upper right region, $a=b\gg c$ and $V\gg c$, the system corresponds to the 1/2 VL phase. The system has almost no type $c$ vertices. At the same time $V$ drives the system to form vortices consisting of vertices of type $a$, $b$ only, as shown in Fig. A2 (a).
@@
In the disordered phase, the characteristics of vortices vary across different parameter regimes. As depicted in Fig. A2 (e), two distinct regimes are identified, labeled as circled-3 and circled-4.
@@
In fact, 1/2 VL is accompanied by a $x$-direction ferromagnetic order and 1/3 VL is accompanied by an AFM order (see appendix A).
@@ IV.2 Vortex structure factors
To further understand the spin vortex lattice phase, the structure factor in $\bm{q}$-space is introduced as
@@
where $n_{i}$, $n_{j}$ represents 1, 0, in the face of the honeycomb lattices, i.e., a triangular lattice. The symbols $\bm{r_{i}}$ and $\bm{r_{j}}$ are the center coordinates of the vortex. In real space, if the density obeys configurations of the form ($101010\cdots$) or ($010101\cdots$), the wave vector corresponding to the maximum value of $S(\bm{q})$ should be located at $\mathbf{q}=(\pi,0)$ [49].
@@
In Fig. 8 (a), for the 2/3 VL phase, $S(q_{x},q_{y})$ is obtained by using Eq. 21 with a lattice size $L=12$. One of the brightest points is located at:
@@
as indicated by the red arrow. The position of the peaks reflects the translational symmetry of the vortex lattice. Assuming that the side length of the honeycomb lattice is 1, the spacing between the two pink vortices $\bm{r_{i}}$ and $\bm{r_{j}}$ is:
@@
The position in Eq. 22 can be obtained by using:
@@
For the 1/2 VL lattice, $S(q_{x},q_{y})$ is shown in Fig. 8 (b). The brightest point (as indicated by the blue arrow) is located at:
@@
and the second brightest point (as indicated by the red arrow) is located at:
@@
This is because in real space, the translation vectors are respectively $\Delta x=2\sqrt{3}$ and $\Delta y=3$.
In total, we have identified two VL phases in the bilayer 6V model with vortex weight where $V>1$. These ordered states can be distinguished by examining either their respective structure factors or the relevant order parameters.
@@ IV.3 Size effects from periodic boundaries
Figure 9 shows the size effects of the vortex density in a 2/3 VL with $L=4,6,8,10$, respectively. The honeycomb lattice is a complex lattice with two sites in its smallest cell, hence having an even number of lattice sites in the horizontal direction, i.e., $L~{}mod~{}2=0$.
@@
Table 3: The table shows the maximum theoretical value of the vortex density $\rho_{v}$ as a function of system size $L$.
@@
where the symbol $\lfloor$.$\rfloor$ . ⌋ denotes rounding down. We first explain Eq. 28 using $L=6$, two sites labeled as 1, 2 in the horizontal direction, a total of $L/2$ bricks in each row as shown in the denominator.
@@
Through the analysis above and the densities presented in Table 3, in the 2/3 VL phase, the density remains constant for sizes where $L~{}mod~{}6=0$. Therefore, when observing physical quantities later on, we only simulate systems whose size is a multiple of 6.
@@ IV.4 Detailed transitions between the phases
IV.4.1 The Ising transition between the AFM and disordered phases at $V<1$
@@
In Figs. 10 (a) and (b), the specific heat $C_{V}$ and $C_{vx}$ exhibit divergence with respect to sizes during the phase transition between the AFM and disordered phases. This behavior stands in contrast to the convergence of $C_{V}$ in the single-layer 6V $F$ model [26], as shown in Fig. 6 (d).
@@
The possible reasons of the Ising transition between the AFM and disordered phases at $V<1$. Be precise are as follows: $V<1$ implies the absence of the disordered phase as depicted by the absence of vortices in Fig. A2 (c). Additionally, the AFM phase has no vertex, as illustrated in Fig. 7 (d). Consequently, the phase transition between these two phases does not involve vortices. Moreover, the transition from simple AFM order to disorder involves $Z2$ symmetry breaking.
Figure 10: The Ising transition between the AFM and disordered phases at $V<1$, $a=b$ and $V=0.5$. (a) $C_{V}$ (b) $C_{vx}$, (c) $Q$ (d) data collapse of $Q$, using $y_{t}=1$. The system size ranges from 8 to 24.
IV.4.2 Transitions between the D and 1/2 VL phases at $V>1$
The phase transition between disordered and the 1/2 VL phase is discussed. By fixing the parameter $V=4$, and scanning $a$ and keeping $a=b$, the vortex density $\rho_{vx}$, and $C_{V}$ are measured for different sizes $L=16-48$.
As shown in Figs. 11 (a) and (b), the jumps of $\rho_{vx}$ and peaks of $C_{V}$ show the signatures of phase transitions. The divergence of $C_{V}$ confirms the phase transition is not of BKT type [30]. In Figs. 11 (c)-(d), the histograms of $E/N$ and $m_{s}$ are obviously double peaked at the parameters $a=b=2$. This indicates the phase transition is of first order.
The signature of the first order transition can also obtained by fitting the critical exponent. Using the package for finite size scaling [50], $\nu$ is obtained as $0.48(8)$ using Eq. 20.
The data for sizes $L=32,36,40$ overlap very well. The scaling dimension $y_{t}=1/0.48=2.08\approx 2$ equals to the system dimension $d=2$ when the first order transition occurs [51].
Figure 11: The signatures of the first order phase transition at $V=4$. The quantities (a) $\rho_{vx}$ (b) $C_{V}$ with system sizes $L=16-48$. The double peaked distributions (c) $P(E/N)$ and (d) $P(m_{s})$ at $a=2.004(1)$. (e) $Q_{s}$ and (f) data collapse of $Q_{s}$ and obtaining $\nu=0.48(8)$ [50].
IV.4.3 Transition between the D and 2/3 VL phases
Figure 12: The signatures of the first order phase transition. The double peaked distributions of (a) $P(E/N)$ and (b) $P(\rho_{vx})$ at $V=7$ and $a=0.35$. (c) $C_{V}$ versus $a$, (d) $a_{c}$ versus $1/L$. (e) $E/N$ versus MC steps and the convergent time is about $10^{6}$.
The transition from the 2/3 VL phase to the disorder phase is also first order. In Figs. 12 (a) and (b), the double peaks in the distribution of $P(E)$ and $P(\rho_{vx})$ indicating a first order phase transition.
Different from the 1/2 VL phase as shown in Fig. 11, this phase transition has an obvious size effect. Figure 12(c) illustrates that the specific heat peak shifts to the right as the size increases. Fortunately, through finite-size scaling defined as follows:
@@ IV.4.3 Transition between the D and 2/3 VL phases
The error bar (2) from 0.475811(2) is calculated by the following equations. First, if one fits $y=kx+b$, then the standard deviation of the intercept b is
@@
where $n$ is the number of points involved, $s_{y}$ is the standard deviation of the observation $y_{i}$ and can be expressed as:
@@
where $n-2$ is the degree of freedom. If one fits the data using “gnuplot” software, the result is $a_{c}\left(\infty\right)=0.476(5)$ consistent with the above result within the error bar.
To show that the results are convergent, Fig. 12 (e) displays $E/N$ versus MC steps. The converging time is about $10^{6}$ steps. The converged energy is approximately 0.12, and denoted as $e_{c}$ which is also marked by a dashed line in Fig. 12 (a), sitting between the two peaks.